OpenStudy (anonymous):

subtract (x-7)/(x-6)-(x-3)/(6-x) a.x-5/x-6 b.2(x-5)/x-6 c.4 d.2(x-4)/x-6

OpenStudy (lalaly):

u need a common denominator \[\frac{x-7}{x-6} - (\frac{x-3}{6-x} \times \frac{-1}{-1}) \]\[=\frac{x-7}{x-6}-\frac{3-x}{x-6} = \frac{x-7-3-(-x)}{x-6}\]\[=\frac{2x-10}{x-6}\]\[=\frac{2(x-5)}{x-6}\]

OpenStudy (anonymous):

THANKSS

OpenStudy (lalaly):

:D:D